# Adapting Large Language Models for Document-Level Machine Translation

### v3

@@ Abstract
Large language models (LLMs) have significantly advanced various natural language processing (NLP) tasks. Recent research indicates that moderately-sized LLMs often outperform larger ones after task-specific fine-tuning. This study focuses on adapting LLMs for document-level machine translation (DocMT) for specific language pairs. We first investigate the impact of prompt strategies on translation performance and then conduct extensive experiments using two fine-tuning methods, three LLM backbones, and 18 translation tasks across nine language pairs. Our results show that specialized models can sometimes surpass GPT-4 in translation performance but still face issues like off-target translation due to error propagation in decoding. We provide an in-depth analysis of these LLMs tailored for DocMT, examining translation errors, discourse phenomena, training strategies, the scaling law of parallel documents, recent test set evaluations, and zero-shot crosslingual transfer. Our findings highlight the strengths and limitations of LLM-based DocMT models and provide a foundation for future research.
@@ 1 Introduction
In this study, we analyze moderately-sized LLMs (with $7B$ parameters) across 18 translation tasks involving nine language pairs. We fine-tune three LLMs using Parameter-Efficient Fine-Tuning (PEFT) and Fully Fine-Tuning (FFT). Comparisons with state-of-the-art translation models, using metrics like $s$BLEU, $d$BLEU, and COMET, confirm the superior translation capabilities of LLMs after fine-tuning. However, we identify a significant issue of off-target translations, observed even after exclusive fine-tuning on bilingual corpora. Our further analysis shows that the high off-target translation rate results from the error propagation in the decoding process. Additionally, we present an in-depth analysis of our LLM-based DocNMT models from various perspectives: translation error distribution, discourse phenomena, training strategy, the scaling law of parallel documents, additional evaluations on WMT2023 test sets (Koehn et al., 2023), and zero-shot cross-lingual transfer, aiming to enhance understanding and efficacy of LLMs in DocMT tasks.
@@ 3 Experimental Setup
In this study, we aim to adapt multilingual pre-trained large language models (LLMs) into a bilingual document-level machine translation (DocMT) model. In this section, we describe our experimental setup of this work, including training strategy (Section 3.1), datasets (Section 3.2), models (Section 3.3), and evaluation (Section 3.4).
@@ 3.4 Evaluation
We evaluate the translation quality using sentence-level BLEU (Papineni et al., 2002) and document-level BLEU (Liu et al., 2020) using SacreBLEU (Post, 2018), denoted as $s$BLEU and $d$BLEU. [^4] BLEU signature: nrefs:1|case:mixed|eff:no|
@@
We use beam search with the beam size of 5 during translation. As shown in 1(d), previous translations serve as the context for the current translation, so the test examples are translated in their original order, beginning with the first sentence free from context.
@@ 4 A Preliminary Study on Prompts
Llama2-7B 02.7 03.5 40.1 04.2 04.4 52.2
Bloom-7B 02.5 02.9 35.5 06.7 07.3 49.4
Vicuna-7B 10.2 12.4 64.7 09.5 09.8 62.9
@@
LLM-based DocMT models (Ours)
@@ Conclusion
Figure 2: Breakdown results on $s$BLEU, $d$BLEU, and COMET given by L-7B-LoRA, V-7B-LoRA, B-7B-LoRA, Doc2Doc-mT5-1.2B, and gpt-4-turbo for the translation tasks from other languages to English.
@@ 5 Main Results
As shown in Table 2, models using LoRA generally outperform fully fine-tuned (FFT) LLMs. However, V-7B-fft surpasses V-7B-LoRA in translating other languages to English, likely due to overfitting. Extensive fine-tuning with a large corpus often leads to rapid overfitting, whereas LoRA, which updates only a few parameters, helps prevent this issue. Additionally, L-7B and V-7B models perform similarly, indicating that further fine-tuning with instruction-tuned models does not always improve task-specific performance.
@@
We present the results for translation tasks from other languages to English in Figure 2. For clarity, only the results from our models using LoRA are shown. Our LLM-based DocMT models generally perform better, sometimes even surpassing gpt-4-turbo in certain tasks, though they fail completely in others. A manual review indicates that the primary cause of these failures is off-target translation. An in-depth analysis of this issue is provided in Section 6, and a complete breakdown of the results is available in Appendix E.
@@ 6 Analysis
In this section, we examine off-target issues with decoding strategy and use GPT-4 to analyze translation errors. We also explore discourse phenomena, training strategies, the scaling law of parallel documents, and additional evaluations on recent WMT2023 test sets. Additionally, we evaluate the zero-shot crosslingual transfer capabilities of our models and present the results in Appendix I.
@@
Decoding Strategy
As described in Section 3.4, previous translations provide context for the current translation. We hypothesize that the high off-target translation rate in Table 3 is due to error propagation during the decoding stage. Therefore, we refer to the decoding strategy in Section 3.4 as ReUse, and introduce an alternative strategy where all context translations are re-generated, called ReGen. As shown in Table 4, the ReGen strategy significantly improves translation quality, confirming our hypothesis on decoding error propagation. However, the inference cost of ReGen is four times higher than that of ReUse. These findings highlight the main reason for translation failures in LLM-based DocMT models, offering insights for future research.
@@
Figure 4: COMET-Percentage (%) of training data for the translations from English to German.
Scaling Law of Parallel Documents
We explore the scaling law for fine-tuning parallel documents. Results for English-German translation are presented in Figure 4, and for English-Romanian and English-Chinese in Appendix H. While LLMs typically excel with minimal training data, different fine-tuning strategies show distinct scaling behaviors. Our LoRA models match full training set performance with just $10\%$ of the data (around $20K$ examples), while fully fine-tuned models achieve near-equivalent performance with only about $1\%$ of the data (approximately $2K$ examples). These are crucial for low-resource languages, as recent LLMs are mainly pre-trained on English text.
@@
Given their extensive pre-training on large text corpora, LLMs are susceptible to data leakage risks. We evaluate our models using recent test sets from WMT2023 (Koehn et al., 2023). These tests, conducted between English and German, assess the out-of-domain generalization of our models and help mitigate data leakage risks. We use spaCy to segment documents and discard any parallel documents where the source and target sides have a differing number of sentences. Our findings, presented in Table 7, reveal that while DocNMT5 models outperform LLM-based models in Table 2, LLM-based models excel in translating out-of-domain text on the WMT2023 test sets. These results highlight the ability of LLM-based DocNMT to generalize well to out-of-domain translation tasks.
@@
$\mu_{\Delta}$ Ar De Fr Zh
L-7B-LoRA +29.4 +36.3 +38.8 +37.2 +29.5
L-7B-fft +29.0 +41.2 +40.5 +37.1 +37.5
B-7B-LoRA +20.3 0+7.5 +40.7 +20.7 +9.8
B-7B-fft +27.3 +14.8 +37.8 +28.9 +19.5
V-7B-LoRA 0-8.9 -12.6 +22.1 +18.9 -34.1
V-7B-fft 0-1.4 0+7.3 +25.2 +17.7 0-3.5
Table 8: The difference ($\Delta$) in COMET scores on the test sets from English to other languages between our English-German LLM-based DocMT models and their backbones. $\mu_{\Delta}$ indicates the average difference across all the languages in this table. The complete results are presented in Appendix I.
Zero-Shot Crosslingual Transfer
@@ 7 Conclusion
This study investigates the adaptation of LLMs for DocMT through extensive experimentation. The research involves two fine-tuning methods, three LLM backbones, and 18 translation tasks across nine language pairs. Our findings indicate that while specialized models sometimes surpass GPT-4 in translation accuracy, they continue to face challenges such as off-target translation due to error propagation during decoding. We present a detailed analysis of these LLMs tailored for DocMT, exploring aspects such as translation errors, discourse phenomena, training strategies, the scaling law of parallel documents, recent test set evaluations, and zero-shot crosslingual transfer. The results highlight both the strengths and limitations of LLM-based DocMT models, offering a foundation for future research in this field.
@@ Appendix D GPT-4 Prompts
Llama2-7B 02.8 00.4 06.6 04.5 01.0 01.0 01.9 00.2 01.9 07.4
Bloom-7B 02.5 01.0 01.0 012.1 01.4 00.1 03.1 00.7 00.1 03.4
Vicuna-7B 10.2 04.5 06.4 06.4 08.6 10.2 09.8 13.9 06.8 25.4
@@
MR-Doc2Sen-mT5 -1.2B 18.8 10.2 18.8 25.6 22.3 14.5 16.2 19.6 19.3 22.8
MR-Doc2Doc-mT5 -1.2B
DocFlat-mT5 -1.2B 19.2 11.0 19.2 25.7 22.6 14.7 16.5 20.3 19.2 23.8
IADA-mT5 -1.2B 19.3 11.7 19.4 26.3 23.9 15.2 16.9 20.9 19.6 23.4
@@
Llama2-7B 03.5 00.5 07.4 04.9 01.1 01.8 03.7 00.2 02.2 09.6
Bloom-7B 02.8 01.0 01.3 12.9 01.7 00.3 02.7 01.0 00.1 04.4
Vicuna-7B 12.4 05.7 06.4 07.1 08.5 15.0 16.1 14.4 07.4 31.2
@@
MR-Doc2Sen-mT5 -1.2B 21.9 11.9 20.7 27.9 23.8 19.8 23.3 21.5 20.7 27.9
MR-Doc2Doc-mT5 -1.2B 22.5 12.1 20.8 28.0 24.7 20.9 24.3 21.9 21.7 27.9
DocFlat-mT5 -1.2B 22.4 12.2 21.3 28.4 24.0 20.8 24.1 21.1 21.4 28.1
IADA-mT5 -1.2B 22.4 12.7 21.7 28.7 24.2 21.4 24.3 21.6 21.7 28.0
@@
Llama2-7B 40.1 37.4 41.5 41.2 39.3 39.7 42.8 35.5 41.6 42.1
Bloom-7B 35.5 34.9 34.8 45.3 33.0 33.9 35.5 34.2 28.6 39.0
Vicuna-7B 64.7 68.3 48.7 49.0 62.5 81.3 72.4 64.1 56.1 80.2
@@
MR-Doc2Sen-mT5 -1.2B 79.9 82.2 76.3 76.5 80.0 83.6 83.1 76.7 79.7 80.9
MR-Doc2Doc-mT5 -1.2B
DocFlat-mT5 -1.2B 80.4 81.8 77.3 76.9 80.3 83.6 83.4 78.0 80.7 81.8
IADA-mT5 -1.2B 80.7 82.4 77.0 77.3 80.9 84.1 83.7 77.8 80.9 81.8
@@
Llama2-7B 04.2 00.1 06.2 01.3 04.4 00.0 00.1 15.1 10.3 00.1
Bloom-7B 06.7 04.7 08.7 14.3 16.9 00.1 00.3 09.4 05.5 00.2
Vicuna-7B 09.5 01.1 14.4 24.9 14.1 04.7 00.4 17.2 08.5 00.0
@@
MR-Doc2Sen-mT5 -1.2B 22.0 26.9 22 29.9 27.7 11.8 13.9 26.5 26.1 18
MR-Doc2Doc-mT5 -1.2B
DocFlat-mT5 -1.2B 22.2 26.6 22.3 29.7 28.4 11.7 13.9 26.4 25.9 17.6
IADA-mT5 -1.2B 22.1 26.9 22.7 30.5 28.5 12.8 14.5 27.1 26.1 18.7
@@
V-7B-fft 21.8 02.2 31.0 43.4 45.0 00.0 00.7 36.2 38.0 00.1
@@
Llama2-7B 04.4 00.1 06.9 01.5 04.7 00.0 00.1 15.7 10.9 00.1
Bloom-7B 07.3 05.3 09.8 015.2 017.6 00.1 00.5 10.6 06.6 00.2
Vicuna-7B 09.8 01.1 14.5 24.6 14.4 05.9 00.5 18.5 08.9 00.0
@@
MR-Doc2Sen-mT5 -1.2B 23.8 27.4 24.2 30.3 29.4 14.9 16.1 27.5 26.8 19.8
MR-Doc2Doc-mT5 -1.2B 24.0 28.3 24.3 30.5 29.8 15.7 16.8 27.8 27.8 20.8
DocFlat-mT5 -1.2B 24.3 27.6 24.5 31.1 29.7 15.1 17.0 28.1 27.8 20.3
IADA-mT5 -1.2B 24.0 28.2 24.6 30.9 29.6 15.0 17.1 27.8 27.1 20.5
@@
Llama2-7B 52.2 50.3 47.1 42.6 51.6 56.1 55.7 56.5 53.0 56.9
Bloom-7B 49.4 50.6 52.8 55.5 56.8 44.5 44.3 51.2 45.2 43.3
Vicuna-7B 62.7 51.3 65.2 70.5 57.3 69.5 56.3 68.0 58.8 67.5
@@
MR-Doc2Sen-mT5 -1.2B 79.8 80.3 79.8 82.3 81.1 76.4 76.6 81.5 81.8 78.2
MR-Doc2Doc-mT5 -1.2B
DocFlat-mT5 -1.2B 80.3 80.2 80.0 82.5 81.7 77.4 77.6 82.2 82.3 78.4
IADA-mT5 -1.2B 80.4 80.3 80.7 82.9 82.3 77.7 77.7 81.8 81.7 78.5
@@
V-7B-fft 74.3 52.4 81.3 87.0 88.2 65.6 55.5 84.2 87.2 67.0

### v4

@@ Abstract
Large language models (LLMs) have significantly advanced various natural language processing (NLP) tasks. Recent research indicates that moderately-sized LLMs often outperform larger ones after task-specific fine-tuning. This study focuses on adapting LLMs for document-level machine translation (DocMT) for specific language pairs. We first investigate the impact of prompt strategies on translation performance and then conduct extensive experiments using two fine-tuning methods, three LLM backbones, and 18 translation tasks across nine language pairs. Our results show that specialized models can sometimes surpass GPT-4 in translation performance but still face issues like off-target translation due to error propagation in decoding. We provide an in-depth analysis of these LLMs tailored for DocMT, examining translation errors, discourse phenomena, strategies for training and inference, the data efficiency of parallel documents, recent test set evaluations, and zero-shot crosslingual transfer. Our findings highlight the strengths and limitations of LLM-based DocMT models and provide a foundation for future research.
@@ 1 Introduction
In this study, we explore the capabilities of moderately-sized Large Language Models (LLMs) with 7 billion parameters across 18 translation tasks involving nine language pairs. We conduct experiments under both bilingual and multilingual settings, employing Parameter-Efficient Fine-Tuning (PEFT) and Fully Fine-Tuning (FFT) techniques on three LLM backbones. Our findings reveal that, after fine-tuning, these LLMs demonstrate superior translation performance compared to state-of-the-art models, as evidenced by metrics such as $s$BLEU, $d$BLEU, and COMET. However, a significant challenge identified is the issue of off-target translations, which persists even after exclusive fine-tuning on bilingual corpora. Our analysis attributes this high rate of off-target translations to error propagation during the inference. Additionally, we provide a comprehensive analysis of our LLM-based DocMT models from various perspectives. This includes examining translation error distribution, discourse phenomena, training strategies, the data efficiency of parallel documents, and additional evaluations on the recent WMT2023 test sets. We also investigate zero-shot cross-lingual transfer, aiming to enhance the understanding and efficacy of LLMs in DocMT tasks.
@@ 3 Experimental Setup
In this section, we detail our experimental setup, covering the training strategy (Section 3.1), datasets (Section 3.2), models (Section 3.3), and evaluation (Section 3.4).
@@ 3.4 Evaluation
We evaluate the translation on the IWSLT2017 test sets (Cettolo et al., 2017) using sentence-level BLEU (Papineni et al., 2002) and document-level BLEU (Liu et al., 2020) using SacreBLEU (Post, 2018), denoted as $s$BLEU and $d$BLEU. [^4] BLEU signature: nrefs:1|case:mixed|eff:no|
@@
Settings
We evaluate our models in two settings: Bilingual English-from/to-Many Translations and Multilingual English-from/to-Many Translations. For the bilingual translations, we train our models on the training set for one specific language pair. For the multilingual translations, we combine all the available training sets for training.
@@
In this work, we explore two inference strategies. The first strategy involves using previous translations as context for the current translation, translating test examples in their original order. This begins with the first sentence, which is free from context, as illustrated in 1(d). We refer to this strategy as ReUse. Alternatively, sentences within the context can be re-generated individually for each document translation to prevent error propagation during the translation process. We call this strategy ReGen. For reasons of inference efficiency, ReUse is used as the default inference strategy throughout this study, unless specified otherwise. A comparison of these two inference strategies is provided in Section 6.
@@ 4 A Preliminary Study on Prompts
Llama2-7B 7B 02.7 03.5 40.1 04.2 04.4 52.2
Bloom-7B 7B 02.5 02.9 35.5 06.7 07.3 49.4
Vicuna-7B 7B 10.2 12.4 64.7 09.5 09.8 62.9
@@
Bilingual English-from/to-Many LLM-based DocMT models (Ours)
@@
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 00.7B 008M 13.9 16.9 67.2 17.1 18.4 62.0
L-7B-fft 00.7B 00.7B 17.1 20.4 73.8 18.3 19.5 69.6
B-7B-LoRA 00.7B 008M 10.3 12.3 58.1 23.5 25.7 75.8
B-7B-fft 00.7B 00.7B 16.0 18.6 67.8 27.2 30.5 78.7
V-7B-LoRA 00.7B 008M 12.9 15.7 66.4 17.2 18.6 60.6
V-7B-fft 00.7B 00.7B 13.8 16.8 69.3 18.4 19.4 65.6
@@ Conclusion
(a) Bilingual English-from-Many translations.
(b) Multilingual English-from-Many translations.
Figure 2: Breakdown results on $s$BLEU, $d$BLEU, and COMET given by L-7B-LoRA, V-7B-LoRA, B-7B-LoRA, Doc2Doc-mT5-1.2B, and gpt-4-turbo for the bilingual (2(a)) and multilingual (2(b)) translation tasks from other languages to English.
@@ 5 Main Results
As shown in Table 2, models using LoRA generally outperform fully fine-tuned (FFT) LLMs in bilingual translations, while fully fine-tuned LLMs surpass LoRA models in multilingual translations. The LoRA approach, which updates only a limited number of parameters, allows the model to retain the rich linguistic knowledge of the pre-trained LLM while adapting its capabilities to a particular bilingual context. However, in multilingual translation tasks, this parameter-efficient fine-tuning is insufficient to capture the complexities of multiple languages simultaneously. Conversely, fully fine-tuning all the parameters of LLMs enables better capture of multilingual complexity but makes the model prone to overfitting in bilingual translations.
@@
We present the results for translation tasks from other languages to English in Figure 2. For clarity, only the results from our models using LoRA are shown. We observe that LLM-based DocMT models sometimes even surpassing gpt-4-turbo in certain bilingual and multilingual translation tasks, though they fail completely in others. Our manual review indicates that the primary cause of these failures is off-target translation. An in-depth analysis of this issue is provided in Section 6, and a complete breakdown of the results is available in Appendix E.
@@ 6 Analysis
In this section, we examine off-target issues with decoding strategy and use GPT-4 to analyze translation errors. We also explore discourse phenomena, training strategies, the data efficiency of parallel documents, and additional evaluations on recent WMT2023 test sets. Additionally, we evaluate the zero-shot crosslingual transfer capabilities of our models and present the results in Appendix I.
@@
Inference Strategy
As described in Section 3.4, previous translations provide context for the current translation. We hypothesize that the high off-target translation rate in Table 3 is due to error propagation during the decoding stage. Therefore, we refer to the decoding strategy in Section 3.4 as ReUse, and introduce an alternative strategy where all context translations are re-generated, called ReGen. As shown in Table 4, the ReGen strategy significantly improves translation quality, confirming our hypothesis on decoding error propagation. However, the inference cost of ReGen is four times higher than that of ReUse. These findings highlight the main reason for translation failures in LLM-based DocMT models.
@@
Figure 4: COMET-Percentage (%) of parallel documents for the translations from English to German.
Data Efficiency of Parallel Documents
We explore the data efficiency of parallel documents. Results for English-German translation are presented in Figure 4, and for English-Romanian and English-Chinese in Appendix H. While LLMs typically excel with minimal training data, different fine-tuning strategies show distinct scaling behaviors. Our LoRA models match full training set performance with just $10\%$ of the data (around $20K$ examples), while fully fine-tuned models achieve near-equivalent performance with only about $1\%$ of the data (approximately $2K$ examples). These are crucial for low-resource languages, as recent LLMs are mainly pre-trained on English text.
@@
Given their extensive pre-training on large text corpora, LLMs are susceptible to data leakage risks. We evaluate our models using recent test sets from WMT2023 (Koehn et al., 2023). These tests, conducted between English and German, assess the out-of-domain generalization of our models and help mitigate data leakage risks. We use spaCy to segment documents and discard any parallel documents where the source and target sides have a differing number of sentences. Our findings, presented in Table 7, reveal that while mT5-based models outperform LLM-based models in Table 2, LLM-based models excel in translating out-of-domain text on the WMT2023 test sets. These results highlight the ability of LLM-based DocMT models to generalize well to out-of-domain translation tasks.
@@ 7 Conclusion
In this study, we adapt large language models (LLMs) for document-level machine translation (DocMT) across specific language pairs and found that moderately-sized, fine-tuned models can sometimes surpass larger models like GPT-4 in translation performance. By investigating prompt strategies and conducting extensive experiments with different fine-tuning methods and LLM backbones, we highlight both the potential and challenges of using LLMs for DocMT. Despite achieving promising results, issues like off-target translations due to error propagation persist. Our analysis of translation errors, discourse phenomena, strategies for training and inference, and data efficiency underscores the strengths and limitations of LLM-based DocMT models, providing a foundation for future research to enhance their effectiveness.
@@ Appendix D GPT-4 Prompts
LLM backbones
Llama2 02.8 00.4 06.6 04.5 01.0 01.0 01.9 00.2 01.9 07.4
Bloom 02.5 01.0 01.0 12.1 01.4 00.1 03.1 00.7 00.1 03.4
Vicuna 10.2 04.5 06.4 06.4 08.6 10.2 09.8 13.9 06.8 25.4
Re-implemented DocMT baselines
@@
MR-Doc2Sen-mT5 18.8 10.2 18.8 25.6 22.3 14.5 16.2 19.6 19.3 22.8
MR-Doc2Doc-mT5
DocFlat-mT5 19.2 11.0 19.2 25.7 22.6 14.7 16.5 20.3 19.2 23.8
IADA-mT5 19.7 11.7 19.4 26.3 23.9 15.2 16.9 20.9 19.6 23.4
Bilingual English-from/to-Many LLM-based DocMT models (Ours)
@@
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 13.9 08.8 22.2 27.9 05.0 07.7 09.4 03.4 19.4 21.6
L-7B-fft 17.1 07.8 22.0 32.5 17.9 12.5 11.5 04.2 20.3 25.1
B-7B-LoRA 10.3 12.5 03.8 36.0 12.8 05.8 01.6 03.7 03.4 12.9
B-7B-fft 16.0 12.6 15.8 37.4 27.7 02.0 01.3 18.4 07.8 21.3
V-7B-LoRA 12.9 08.0 15.4 28.2 06.2 10.1 07.1 03.3 15.1 22.4
V-7B-fft 13.8 10.4 15.2 11.3 08.4 11.1 12.2 12.4 17.0 26.3
@@
LLM backbones
Llama2 03.5 00.5 07.4 04.9 01.1 01.8 03.7 00.2 02.2 09.6
Bloom 02.8 01.0 01.3 12.9 01.7 00.3 02.7 01.0 00.1 04.4
Vicuna 12.4 05.7 06.4 07.1 08.5 15.0 16.1 14.4 07.4 31.2
Re-implemented DocMT baselines
@@
MR-Doc2Sen-mT5
MR-Doc2Doc-mT5 21.9 11.9 20.7 27.9 23.8 19.8 23.3 21.5 20.7 27.9
DocFlat-mT5 22.5 12.1 20.8 28.0 24.7 20.9 24.3 21.9 21.7 27.9
IADA-mT5 22.4 12.2 21.3 28.4 24.0 20.8 24.1 21.1 21.4 28.1
Bilingual English-from/to-Many LLM-based DocMT models (Ours)
@@
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 16.9 10.2 26.0 30.4 05.6 12.8 15.7 03.8 20.2 27.6
L-7B-fft 20.4 09.2 25.5 35.5 18.5 19.2 18.3 04.9 20.9 31.6
B-7B-LoRA 12.3 14.3 04.4 39.5 14.4 10.8 03.3 03.9 03.5 16.8
B-7B-fft 18.6 14.2 19.5 40.9 30.9 03.8 02.6 20.9 08.5 26.5
V-7B-LoRA 15.7 09.5 18.1 31.0 06.5 16.0 11.8 03.9 16.1 28.1
V-7B-fft 16.8 12.0 17.5 12.9 08.8 17.4 19.5 12.4 17.8 32.7
@@
LLM backbones
Llama2 40.1 37.4 41.5 41.2 39.3 39.7 42.8 35.5 41.6 42.1
Bloom 35.5 34.9 34.8 45.3 33.0 33.9 35.5 34.2 28.6 39.0
Vicuna 64.7 68.3 48.7 49.0 62.5 81.3 72.4 64.1 56.1 80.2
Re-implemented DocMT baselines
@@
MR-Doc2Sen-mT5 79.9 82.2 76.3 76.5 80.0 83.6 83.1 76.7 79.7 80.9
MR-Doc2Doc-mT5
DocFlat-mT5 80.4 81.8 77.3 76.9 80.3 83.6 83.4 78.0 80.7 81.8
IADA-mT5 80.7 82.4 77.0 77.3 80.9 84.1 83.7 77.8 80.9 81.8
Bilingual English-from/to-Many LLM-based DocMT models (Ours)
@@
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 67.2 78.8 74.1 69.3 40.9 65.9 74.7 44.9 79.3 77.1
L-7B-fft 73.8 78.8 76.3 77.0 65.2 80.6 79.0 45.3 81.9 79.8
B-7B-LoRA 58.1 80.9 43.8 79.9 54.9 66.7 46.6 38.1 46.9 65.3
B-7B-fft 67.8 82.2 69.6 83.8 81.7 46.3 45.7 70.5 54.5 76.0
V-7B-LoRA 66.4 78.2 65.5 70.8 46.9 74.5 66.4 41.9 76.0 77.6
V-7B-fft 69.3 80.3 64.7 54.2 52.0 78.1 77.5 57.6 78.1 81.1
@@
LLM backbones
Llama2 04.2 00.1 06.2 01.3 04.4 00.0 00.1 15.1 10.3 00.1
Bloom 06.7 04.7 08.7 14.3 16.9 00.1 00.3 09.4 05.5 00.2
Vicuna 09.5 01.1 14.4 24.9 14.1 04.7 00.4 17.2 08.5 00.0
Re-implemented DocMT baselines
@@
MR-Doc2Sen-mT5 22.5 26.9 22.0 29.9 27.7 11.8 13.9 26.5 26.1 18.0
MR-Doc2Doc-mT5
DocFlat-mT5 22.5 26.6 22.3 29.7 28.4 11.7 13.9 26.4 25.9 17.6
IADA-mT5 23.1 26.9 22.7 30.5 28.5 12.8 14.5 27.1 26.1 18.7
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
@@
V-7B-fft 24.3 24.2 31.0 43.4 45.0 00.0 00.7 36.2 38.0 00.1
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 17.1 02.3 15.7 18.6 41.6 00.4 00.8 37.6 36.7 00.3
L-7B-fft 18.3 15.8 10.3 22.0 40.7 02.1 06.0 29.7 30.8 07.8
B-7B-LoRA 23.5 34.4 28.5 39.5 40.0 00.1 00.6 33.7 34.1 00.6
B-7B-fft 27.2 35.9 28.3 39.7 38.8 07.1 11.2 32.4 33.5 17.9
V-7B-LoRA 17.2 06.1 05.5 12.2 43.3 02.9 01.8 35.9 36.4 10.2
V-7B-fft 18.4 13.0 17.7 16.7 37.6 02.5 03.7 36.2 35.2 03.0
@@
LLM backbones
Llama2 04.4 00.1 06.9 01.5 04.7 00.0 00.1 15.7 10.9 00.1
Bloom 07.3 05.3 09.8 15.2 17.6 00.1 00.5 10.6 06.6 00.2
Vicuna 09.8 01.1 14.5 24.6 14.4 05.9 00.5 18.5 08.9 00.0
Re-implemented DocMT baselines
@@
MR-Doc2Sen-mT5 24.0 27.4 24.2 30.3 29.4 14.9 16.1 27.5 26.8 19.8
MR-Doc2Doc-mT5 24.6 28.3 24.3 30.5 29.8 15.7 16.8 27.8 27.8 20.8
DocFlat-mT5 24.6 27.6 24.5 31.1 29.7 15.1 17.0 28.1 27.8 20.3
IADA-mT5 24.5 28.2 24.6 30.9 29.6 15.0 17.1 27.8 27.1 20.5
Bilingual English-from/to-Many LLM-based DocMT models (Ours)
@@
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 18.4 02.2 17.6 19.5 44.6 00.4 00.8 40.4 39.7 00.2
L-7B-fft 19.5 15.4 11.1 23.3 43.5 02.4 07.0 31.4 33.0 08.5
B-7B-LoRA 25.7 37.9 32.1 42.5 43.7 00.2 00.7 36.7 37.5 00.5
B-7B-fft 30.5 39.5 31.9 42.6 41.9 10.0 16.0 35.3 37.0 20.6
V-7B-LoRA 18.6 06.4 06.0 13.5 46.6 03.5 02.1 38.3 39.7 11.6
V-7B-fft 19.4 13.4 17.1 17.4 39.6 02.9 04.4 37.8 38.1 03.6
@@
LLM backbones
Llama2 52.2 50.3 47.1 42.6 51.6 56.1 55.7 56.5 53.0 56.9
Bloom 49.4 50.6 52.8 55.5 56.8 44.5 44.3 51.2 45.2 43.3
Vicuna 62.7 51.3 65.2 70.5 57.3 69.5 56.3 68.0 58.8 67.5
Re-implemented DocMT baselines
@@
MR-Doc2Sen-mT5 79.8 80.3 79.8 82.3 81.1 76.4 76.6 81.5 81.8 78.2
MR-Doc2Doc-mT5
DocFlat-mT5 80.3 80.2 80.0 82.5 81.7 77.4 77.6 82.2 82.3 78.4
IADA-mT5 80.4 80.3 80.7 82.9 82.3 77.7 77.7 81.8 81.7 78.5
Bilingual English-from/to-Many LLM-based DocMT models (Ours)
@@
V-7B-fft 77.0 77.0 81.3 87.0 88.2 65.6 55.5 84.2 87.2 67.0
Multilingual English-from/to-Many LLM-based DocMT models (Ours)
L-7B-LoRA 62.0 58.9 54.0 63.8 83.7 42.5 54.5 73.7 77.2 50.0
L-7B-fft 69.6 52.9 63.2 63.2 85.1 65.3 57.0 86.3 86.3 66.8
B-7B-LoRA 78.7 84.0 82.5 86.5 84.6 62.4 68.4 82.3 84.2 73.8
B-7B-fft 75.8 81.0 83.3 86.2 86.9 59.1 51.7 83.9 85.7 64.4
V-7B-LoRA 60.6 52.4 58.2 57.3 78.8 43.6 47.6 83.3 83.8 40.6
V-7B-fft 65.6 50.9 47.4 53.6 87.2 61.4 56.8 83.3 86.3 63.5